I need to write these two problems in the following in Ax+By=C Form
1.) with slope 2/3 through (-1,-4)

2.) through (-4,5) and (-2,-3)

Answer by stanbon(75887)

write these two problems in the following in Ax+By=C Form
1.) with slope 2/3 through (-1,-4)
Form: y = mx + b
You know y=-4 when x=-1 and you know m=2/3; solve for "b":
-4 = (2/3)(-1) + b
-4 = -2/3 + b
b = -10/3
EQUATION:
y = (2/3)x - (10/3)
-----------------------------
2.) through (-4,5) and (-2,-3)
slope = (-3-5)/(-2--4) = -8/2 = -4
Form: y = mx + b
-3 = -4*-2 + b
-3 = 8 + b
b = -11
------------
y = -4x - 11
==================
Cheers,
Stan H.